OpenStudy (anonymous):

Given that the acceleration of gravity at the surface of Mars is 0.38 of what it is on Earth, and that Mars' radius is 3400 km, determine the mass of Mars.

OpenStudy (kmeis002):

Do you know the equation for local gravitational acceleration?

OpenStudy (anonymous):

No not sure, but could it be g= GM/R^2

OpenStudy (kmeis002):

Yes, and we know that g of earth is ~ 9.81, so mars is .38(g_earth). You should now know every thing except M using that knowledge.

OpenStudy (anonymous):

using that above i got a very large number as in: 298430526.3

OpenStudy (kmeis002):

What is the mass refering to? Is it fine to expect a large number for the mass of mars? What units are on this number?

OpenStudy (anonymous):

i am assuming so, and kg

OpenStudy (anonymous):

9.81=(0.38)M/(3400)^2 Right?

OpenStudy (kmeis002):

Ah, Not quite, G is the gravitational constant, it should be .38*9.81 = GM/(3400)^2. You can look up G.

OpenStudy (loser66):

is it not \[\large g_{mars} = G\frac{m_{mars}}{r_{mars}^2}\] where \(g_{mars}= 0.38 g_{earth}= 0.38 *9.8 = 3.424\) so, \[\large m_{mars}= \frac{3.424 * r_{mars}^2}{G}\]

OpenStudy (loser66):

sorry, 3.724 not 3.424

OpenStudy (loser66):

I think I am right , I checked on internet my answer, it's correct XD

OpenStudy (anonymous):

i am getting..6.451287277*10^17

OpenStudy (loser66):